OpenStudy (anonymous):

simplify -6x2-8y3+2x2-5y3 a. -4x2+13y3 b.-4x2-13y3 c.-4x2+12y3 d.4x2-13y3

OpenStudy (redohawk):

B

OpenStudy (anonymous):

add the terms which have the same symbol -4 -13

OpenStudy (salazarblack):

didn't you already post this one?

OpenStudy (redohawk):

-6x2 + 2x2 = -4x2 -8y3 - 5y3 = -13y3 so final = -4x2-13y3

OpenStudy (imtiaz7):

b is correct

OpenStudy (anonymous):

thanks y'all

OpenStudy (anonymous):

@Raven_Draeving You are posting questions with the same idea. it would be better if you understand the concept and do the others on your own.

OpenStudy (redohawk):

NP, happy to help ^^

OpenStudy (camerondoherty):

Step 1 : Simplify \[-4x^2-8y^3 - 5y^3\] Pull out like terms: \[ -4x^2 - 13y^3 = -1 • (4x^2 + 13y^3) \] Factor: 4x2 + 13y3 Theory : A sum of two perfect cubes, \[a^3 + b^3\] can be factored into : \[ (a+b) • (a^2-ab+b^2)\] Check: \[(a+b) • (a^2-ab+b^2)\] \[ a^3-a^2b+ab^2+ba^2-b^2a+b^3 \] \[a^3+(a^2b-ba^2)+(ab^2-b^2a)+b^3\] \[ a^3+0+0+b^3=\] \[ a^3+b^3\] Check : 4 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Answer : \[ -1 • (4x^2 + 13y^3)\] Which Simplifies to: \[-4x^2-13y^3\]